# Higgs Mechanism for the Ashtekar Self-Dual Connection

### v1

@@ Abstract
We introduce the Higgs mechanism for the self-dual spin connection (also known as the Ashtekar connection), using the Plebański formulation of gravity. We develop our formalism within the framework of the chiral action and derive the equations of motion of the theory. One particular test model is explored: since anisotropy is an intrinsic property of the theory, a modified version of the spatially flat Bianchi I model with two different scale factors is considered. We apply our formalism and derive the Friedmann equations which regulate the scale factors and the Higgs field. We also present a Proca-like term for the connection, which when reduced to minisuperspace with a positive $\Lambda$ yields a De Sitter universe with the solution scaled by a constant that depends on the mass of the gauge fields. We finally investigate the effect of these mass terms on gravitational waves and find that the wave equation remains unchanged relatively to GR; however the Weyl tensor is scaled by a constant which depends on the mass of the connection components.
@@ III.1 Modified Bianchi I Model
Equations (49) and (51) correspond to the anti-self-dual part and yield
$\frac{\ddot{b}}{b}=(1+\ell_{P}^{2}v^{2})\left(\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}\right)-\frac{\dot{a}\dot{b}}{ab}(1-\ell_{P}^{2}v^{2}).$ (52)
On the other hand equations (48) and (50) correspond to the self-dual part and are under the constraint ${\rm tr}\,M=\ell_{P}^{2}V$, which originates
$\left(\frac{\ddot{a}}{a}+\frac{\dot{a}\dot{b}}{ab}\right)(1-\ell_{P}^{2}v^{2}/2)+$
$+\frac{\dot{a}^{2}}{2a^{2}}+\frac{\ddot{b}}{2b}+\frac{\ell_{P}^{2}\dot{\varphi}_{3}^{2}}{4}=\ell_{P}^{2}V.$ (53)
The remaining equation to consider is the equation of motion for the scalar field (24) in this spacetime, which takes the form:
@@ IV Proca Term for the Ashtekar Connection
$\frac{\dot{a}^{2}}{a^{2}}=\frac{2+m^{2}\ell_{P}^{2}}{2}\frac{\Lambda}{3}.$ (63)
This is nothing but the usual solution for FLRW in GR with a cosmological constant $\Lambda$ (de Sitter (dS) space, $\Lambda>0$) but with an extra constant that depends on the mass of the gauge fields.
@@ VI Conclusions and Outlook
We also considered a Proca-like term for the Ashtekar connection and therefore all its components became massive. When reduced to MSS with a cosmological constant $\Lambda$ this led to the usual dS universe solution with an extra constant which depends on the mass of the connection. Moreover, we applied this model to gravitational waves with tensorial perturbations in the Plebański formalism and we found that for a Minkowski background the gravitational wave equation stays unchanged and only the Weyl tensor gets scaled by a constant also dependent on the mass of the connection.

### v2

@@ Abstract
We introduce the Higgs mechanism for the self-dual spin connection (also known as the Ashtekar connection), using the Plebański formulation of gravity. We develop our formalism within the framework of the chiral action and derive the equations of motion of the theory. One particular test model is explored: since anisotropy is an intrinsic property of the theory, a modified version of the spatially flat Bianchi I model with two different scale factors is considered. We apply our formalism and derive the Friedmann equations which regulate the scale factors and the Higgs field. We also present a Proca-like term for the connection, which when reduced to minisuperspace with a positive $\Lambda$ yields a De Sitter universe with an effective cosmological constant that depends on the mass of the gauge fields. We finally investigate the effect of these mass terms on gravitational waves and find that the wave equation remains unchanged relatively to GR; however the Weyl tensor is scaled by a constant which depends on the mass of the connection components.
@@ III.1 Modified Bianchi I Model
The first Friedmann equation for this universe can be obtained by combining (48)--(49) and (50)--(51) together with the constraint ${\rm tr}\,M=\ell_{P}^{2}V$:
$2\frac{\dot{a}\dot{b}}{ab}+\frac{\dot{a}^{2}}{a^{2}}(1-\ell_{P}^{2}v^{2})=\ell_{P}^{2}\left(\frac{\dot{\varphi}_{3}^{2}}{2}+V\right).$ (52)
For the second Friedmann equation we do (48)$+$(49) and (50)$+$(51) together with the constraint ${\rm tr}\,M=\ell_{P}^{2}V$:
$2\frac{\ddot{a}}{a}+\frac{\ddot{b}}{b}-\ell_{P}^{2}v^{2}\left(\frac{\ddot{a}}{a}+\frac{\dot{a}\dot{b}}{ab}-\frac{\dot{a}^{2}}{a^{2}}\right)=\ell_{P}^{2}\left(-\dot{\varphi}_{3}^{2}+V\right).$ (53)
One can see that setting $v=0$ recovers the usual GR Friedmann equations. The remaining equation to consider is the equation of motion for the scalar field (24) in this spacetime, which takes the form:
@@ IV Proca Term for the Ashtekar Connection
$\frac{\dot{a}^{2}}{a^{2}}=\frac{\Lambda_{\text{eff}}}{3}.$ (63)
This is nothing but the usual equation for the scale factor in FLRW spacetime in GR but with an effective cosmological constant given by
$\Lambda_{\text{eff}}=\frac{2+m^{2}\ell_{P}^{2}}{2}\Lambda.$ (64)
The solution then depends on the mass of the gauge fields and on $\Lambda$. We have $\Lambda_{\text{eff}}>0$ and therefore a real scale factor, representing a dS spacetime.
@@ VI Conclusions and Outlook
We also considered a Proca-like term for the Ashtekar connection and therefore all its components became massive. When reduced to MSS with a cosmological constant $\Lambda$ this led to the usual dS universe solution with an effective cosmological constant which depends on the mass of the connection. Moreover, we applied this model to gravitational waves with tensorial perturbations in the Plebański formalism and we found that for a Minkowski background the gravitational wave equation stays unchanged and only the Weyl tensor gets scaled by a constant also dependent on the mass of the connection.